# Distributed Motion Control of Multiple Mobile Manipulators for Reducing Interaction Wrench in Object Manipulation

### v2

@@ Abstract
In real-world cooperative manipulation of objects, multiple mobile manipulator systems may suffer from disturbances and asynchrony, leading to excessive interaction wrenches and potentially causing object damage or emergency stops. This paper presents a novel distributed motion control approach aimed at reducing these unnecessary interaction wrenches. The control strategy for each robot only utilizes information from the local force sensors and neighboring robots, without the need for global position and velocity information. Disturbances are corrected through compensatory movements of the manipulators. Besides, the robustness of the control law against communication delays between robots is also considered. The stability of the control law is rigorously proven by the Lyapunov theorem. Subsequently, the efficacy of the proposed control law is validated through simulations and experiments of collaborative object manipulation by two robots. Experimental results demonstrate the effectiveness of the proposed control law in reducing interaction wrenches during object manipulation.
Note to Practitioners
This work is motivated by reducing the interaction wrenches in the cooperative manipulation of objects by multiple mobile manipulator robots. During manipulation, robots often grip the object tightly, and any disturbance in the motions can result in increased interaction wrenches. These interaction wrenches do not contribute to supporting or manipulating the object but may damage both the object and the robots. Previous methods typically rely on extensive global position and velocity information or torque control of robots, which are not common in industrial settings and robots. Thus, their methods are challenging to validate or apply practically. This work introduces a novel distributed control law for reducing interaction wrenches, which only requires information from the local force sensors and neighboring robots. Moreover, the robustness of the control law against communication delays is considered. The proposed control law is convenient to apply and has been verified in physical robots, demonstrating its practical applicability.
Index Terms:
@@ I Introduction
Mobile manipulator robots combine the extensive mobility of a mobile platform with the dexterous manipulation capabilities of a robotic manipulator. This integration allows them to be used in various industrial applications, including object delivery, assembly, machining, and transportation [1, 2, 3]. As the complexity of industrial tasks rises, the limitations of single robots become apparent. Tasks such as manipulating large objects and performing complex assemblies often require the collaboration of multiple robots [4][5]. Moreover, multi-robot collaboration offers more advantages, such as efficiency, robustness, adaptivity, and flexibility [6]. Therefore, in industrial automation, Multiple Mobile Manipulator Systems (MMMS) have attracted widespread attention in object manipulation. A case is shown in Fig. 1. In MMMS collaboration, establishing desired consistency or formations is crucial, such as coordinated formations, motions, and wrenches. In related research, some studies focus on motion planning problems, mainly including redundant and coordinated planning [7].
@@
However, when executing these planned trajectories in real-world scenarios, achieving perfect consistency becomes impossible due to numerous factors such as interactions with objects and some uncertainties. Due to the tight coupling between robots and the object, even small disturbances may cause excessive interaction wrenches. These unnecessary interaction wrenches do not contribute to supporting the object, instead, they may lead to emergency stops of robots or damage to the object [8][9]. While attentive planning is essential for guiding the robots, robust control strategies that can adapt to uncertainties are equally or even more crucial for achieving reliable performance in practical applications [10].
Cooperative control of MMMS in object manipulation, therefore, is an essential area of research. Based on adaptive impedance control and the leader-follower method, the prescribed performance of MMMS in object transportation was achieved in [11]. To track the reference trajectory and avoid obstacles, a method integrating formation control with constrained optimization for MMMS was proposed in [12]. The constrained optimization was solved by a discontinuous projected primal-dual algorithm with damping terms. A constrained optimization method for multi-robot formation control in dynamic environments was proposed in [13], optimizing formation parameters to avoid collisions and progress toward goals. The approach included local motion planning via sequential convex programming and global path planning through sampling and graph search. The above studies enable MMMS to achieve some desired coordination performance.
Generally, the control strategies of MMMS can be categorized into centralized, leader-follower, and fully distributed methods. Centralized methods employ a central controller to control the movements and motions of all mobile manipulators [14]. These methods benefit from a global perspective, enabling optimal task allocation and coordination strategies. However, centralized methods suffer from scalability issues as the number of robots increases [15]. The computational burden on the central controller grows, and the system becomes vulnerable to a single point of failure. In recent years, more attention has been given to the latter two methods in research.
The leader-follower method simplifies the coordination problem by designating one or more robots as the leaders and the others as followers. The leaders make the primary decisions, and the followers adjust their motions based on the leaders’ state to reduce unnecessary interaction wrenches. In [16], a leader-follower structure was used for MMMS to cooperatively transport objects in a constrained workspace. It combined novel estimation laws and adaptive control to enhance robust motion control and collision avoidance. Leader-follower control of MMMS for collective object transportation was investigated in [17], optimizing for energy efficiency and manipulability. The stability of the closed-loop system was proven, and simulations demonstrated the effectiveness of this optimization scheme. In [18], the control algorithm simulated impedance dynamics at each manipulator’s grasping point, enabling coordinated handling based on a leader-follower approach. In [19] and [20], every follower synchronized to the direction of the force applied by one leader robot through the unified or proportional force-amplifying framework. The leader-follower method was validated in physical experiments in [19, 20, 18]. It reduces the complexity of control for the followers and is easier to implement on real-world robots. However, the system’s overall performance heavily relies on the leader’s reliability and decision-making capabilities, introducing a potential single point of failure similar to centralized methods.
Fully distributed methods represent the most advanced approach, emphasizing decentralized control where each mobile manipulator makes decisions based on local information and communication with neighboring robots. In [21], a distributed adaptive controller was proposed to achieve synchronization based on the desired load distribution and desired internal force. The effectiveness of their method was validated through experiments. A distributed model predictive control method was presented for MMMS to transport objects with respect to state, control input, and synchronization constraints in [22]. The method simplifies the problem into two independent synchronization tracking control issues for task-space end-effectors and null-space mobile bases. In [23], a chatterless integral sliding mode force-position control was proposed. Each robot was designed to correct collectively the velocity deviation due to the pushing of the other robots. Distributed coordination and cooperation control for networked mobile manipulators over a topology was introduced in [24]. Synchronization control combined with force controls ensured cooperative manipulation and transportation. In [25], a control approach was proposed to guarantee predefined transient and steady-state performance for object trajectory. Load distribution between robots was also considered by using a grasp matrix pseudoinverse.
While fully distributed methods have significantly advanced the control and coordination of multiple mobile manipulators, they exhibit certain limitations when applied to real-world scenarios. First, existing methods often depend on accurate global positions or velocities of all robots to achieve effective coordination [21, 24, 23, 25, 26, 12]. In industrial automation environments, obtaining precise global positioning can be very challenging. The reliance on global data makes these control methods less robust and difficult to apply. Second, many of these approaches require torque control of the robots [21, 24, 27, 25, 22]. Torque control demands sophisticated hardware and software capabilities, including high-fidelity sensors which may not be present in all industrial robots. This limitation restricts the applicability of these methods to only those systems that are specifically designed and equipped for torque control. Third, distributed methods often require communication between robots to share local information and maintain coordination. In practice, communication delays are inevitable. The aforementioned methods have not been validated for robustness against such delays.
Therefore, to address these shortcomings and make MMMS more convenient and applicable in real-world object manipulation, this paper presents a distributed motion control strategy. The strategy aims to reduce unnecessary interaction wrenches and does not rely on global position, velocity data, or torque control. Instead, each robot leverages information from the local force sensor and other robots to achieve disturbance correction through manipulator motions, addressing uncertainties such as kinematic inaccuracies. The stability of the proposed control law is demonstrated through the Lyapunov theorem. Additionally, the study explores the relationship between delay limits, control parameters, and object stiffness. Finally, the control law is validated through simulations and physical experiments. Both results show that the method can effectively reduce interaction wrenches.
The main contributions of this paper are twofold. First, the proposed distributed motion control law utilizes local force sensors and inter-robot communication to reduce interaction wrenches without requiring extensive global data or torque control. In contrast, existing methods often rely on these, which can be challenging to obtain in industrial settings. Second, the stability of the proposed control law against communication delays is rigorously proven using the Lyapunov theorem, ensuring robust performance. Then, the control law is implemented on physical robots and the results demonstrate the effectiveness in reducing interaction wrenches.
The remainder of this paper is organized as follows. In Section II, some basic concepts are defined. In Section III, the problem is formulated and the proposed control law is presented. Also, the method is proven theoretically and verified by simulations. In Section IV, experiments are conducted on physical robots and results are analyzed. Finally, Section V concludes the paper and outlines future work.
II Preliminaries
II-A Kinematics of Mobile Manipulator
@@
A mobile manipulator robot usually consists of a mobile platform and a mounted robotic manipulator. This paper considers an omnidirectional mobile platform and a six-degree-of-freedom manipulator. As shown in Fig. 1, three mobile manipulators cooperatively manipulate objects. $\mathbb{W}$ denotes the frame of the world. $\mathbb{P}_{i}$, $\mathbb{B}_{i}$, and $\mathbb{E}_{i}$ denote the frames of the mobile platform, the base of the manipulator, and the end-effector of the $i$-th mobile manipulator, respectively, $i=1,\cdots,n$. Let $\bf{T}$, $\bf{R}$, $\bf{P}$ represent the relationship between different frames. For example, ${{}^{\mathbb{W}}_{\mathbb{E}_{i}}}{\bf{T}}\in SE(3)$ represents the pose of frame $\mathbb{E}_{i}$ in frame $\mathbb{W}$, ${{}^{\mathbb{W}}_{\mathbb{E}_{i}}}{\bf{R}}\in SO(3)$ represents the orientation of frame $\mathbb{E}_{i}$ in frame $\mathbb{W}$, and ${{}^{\mathbb{W}}_{\mathbb{E}_{i}}}{\bf{P}}\in\mathbb{R}^{3}$ represents the translation of frame $\mathbb{E}_{i}$ in frame $\mathbb{W}$. The forward kinematics of the robot is
@@
where $\bm{e}_{i}\in\mathbb{R}^{6}$ is the pose of the end-effector in frame $\mathbb{W}$, $f_{\text{mm}}(\cdot)$ is related to ${{}^{\mathbb{W}}_{\mathbb{E}_{i}}}{\bf{T}}$, and $\bm{q}_{\text{mm},i}\in\mathbb{R}^{9}$ is the configuration of the robot. Specifically, $\bm{q}_{\text{mm},i}=[\bm{q}_{\text{p},i}^{\text{T}},\bm{q}_{i}^{\text{T}}]^{\text{T}}$, where $\bm{q}_{\text{p},i}\in\mathbb{R}^{3}$ is the configuration the mobile platform and $\bm{q}_{i}\in\mathbb{R}^{6}$ is the configuration the manipulator. Differentiating Eq.(1), the velocity relationship is
$\dot{\bm{e}}_{i}=\bm{J}_{\text{mm},i}(\bm{q}_{\text{mm},i})\dot{\bm{q}}_{\text{mm},i}$ (2)
where $\bm{J}_{\text{mm},i}=[\bm{J}_{\text{p},i},\bm{J}_{i}]^{\text{T}}$ is the Jacobian matrix of the robot.
II-B Force Sensor
In order to manipulate objects, the end of the manipulator is usually equipped with a gripper and a force sensor. The data measured by force sensors are usually in their own coordinate systems. Let ${}^{\mathbb{E}_{i}}\bm{h}_{i}=[{{}^{\mathbb{E}_{i}}}\bm{f}_{i}^{\text{T}},{{}^{\mathbb{E}_{i}}}\bm{t}_{i}^{\text{T}}]^{\text{T}}$ represent the data measured by the sensor at the end effector of the $i$-th robot. $\bm{h}_{i}=[\bm{f}_{i}^{\text{T}},\bm{t}_{i}^{\text{T}}]^{\text{T}}$ represents the force and torque exerted on the end of the $i$-th robot in frame $\mathbb{W}$. According to the force transfer relationship, it can be derived
@@ II-B Force Sensor
Before measuring force data, it is necessary to perform calibration and gravity compensation for the force sensor. In unloaded conditions, the force data consists of
@@
where $\bm{f}_{\text{s}\_0}$ and $\bm{t}_{\text{s}\_0}$ are the drift errors, $\bm{f}_{\text{s}\_g}$ and $\bm{t}_{\text{s}\_g}$ are caused by the gravity of the gripper $mg$. The displacement from the sensor to the gripper’s center of mass is denoted as ${{}^{\mathbb{E}}_{g}}{\bf{P}}$. Parameters $\bm{f}_{\text{s}\_0}$, $\bm{t}_{\text{s}\_0}$, ${{}^{\mathbb{E}}_{g}}{\bf{P}}$, and $mg$, which need further calibration, are linearly independent. Therefore, these parameters can be calculated using the least squares method after obtaining enough sensor data in unloaded conditions [28].
@@
where $\bm{G}_{6\times 6n}$ is the grasp matrix, which is only related to the relative poses of the grasping points and the object’s center of mass [29, 30, 21], as the $\bm{r}_{1},\bm{r}_{2},\bm{r}_{3}$ shown in Fig. 1.
II-C Graph Theory
The communication structure of a system with $n$ robots can be typically represented on graphs. Let $\mathcal{G=(V,E)}$ denote a graph, where $\mathcal{V}=\left\{v_{i}\right\}$ is the vertex set and =$\mathcal{E}=\left\{a_{ij}\right\}$ is the edge set. $a_{ij}=1$ if the $i$-th robot can received information from the $j$-th robot, otherwise $a_{ij}=0$. The adjacency matrix of the graph is $A=(a_{ij})\in\mathbb{R}^{n\times n}$. The degree matrix of the graph is $D=\text{diag}\left\{\sum_{j=1}^{n}a_{ij}\right\}\in\mathbb{R}^{n\times n}$. The Laplacian matrix of the graph is $L=D-A$. In this paper, the directed graph is considered, which denotes $a_{ij}\neq a_{ij}$.
III The Proposed Control Method
III-A Problem Statement and Control Law

### v3

@@ Abstract
In real-world cooperative manipulation of objects, multiple mobile manipulator systems may suffer from disturbances and asynchrony, leading to excessive interaction wrenches and potentially causing object damage or emergency stops. Existing methods often rely on torque control and dynamic models, which are uncommon in many industrial robots and settings. Additionally, dynamic models often neglect joint friction forces and are not accurate. These methods are challenging to implement and validate in physical systems. To address the problems, this paper presents a novel distributed motion control approach aimed at reducing these unnecessary interaction wrenches. The control law is only based on local information and joint velocity control to enhance practical applicability. The communication delays within the distributed architecture are considered. The stability of the control law is rigorously proven by the Lyapunov theorem. In the simulations, the effectiveness is shown, and the impact of communication graph connectivity and communication delays has been studied. A comparison with other methods shows the advantages of the proposed control law in terms of convergence speed and robustness. Finally, the control law has been validated in physical experiments. It does not require dynamic modeling or torque control, and thus is more user-friendly for physical robots.
keywords:
@@ 1 Introduction
Mobile manipulator robots combine the extensive mobility of a mobile platform with the dexterous manipulation capabilities of a robotic manipulator. This integration allows them to be used in various industrial applications, including object delivery, assembly, machining, and transportation [1, 2, 3]. As the complexity of industrial tasks rises, the limitations of single robots become apparent. Tasks such as manipulating large objects and performing complex assemblies often require the collaboration of multiple robots [4, 5]. Moreover, multi-robot collaboration offers more advantages, such as efficiency, robustness, adaptivity, and flexibility [6]. Therefore, in industrial automation, Multiple Mobile Manipulator Systems (MMMS) have attracted widespread attention in object manipulation.
In related research, some studies focus on motion planning problems, mainly including redundant and coordinated planning [7]. However, in practice, it is impossible to perfectly achieve the planned results in robot motions due to various real-world factors. These factors include disturbances and asynchrony between the robots, which may result from communication delays or kinematic uncertainties. It can lead to deviations from the planned cooperative performance. Thus, some research aims at cooperative control of MMMS to ensure the desired performance [8]. To track the reference trajectory and avoid obstacles, a method integrating formation control with constrained optimization for MMMS was proposed in [9]. A constrained optimization method for multi-robot formation control in dynamic environments was proposed in [10], optimizing formation parameters to avoid collisions and progress toward goals. The approach included local motion planning via sequential convex programming and global path planning through sampling and graph search. The above studies enable MMMS to achieve some desired coordination performance.
In object manipulation, the end-effector of the robot is usually a gripper and grasps the object. When multiple robots grasp an object simultaneously, it necessitates precise coordination of interaction wrenches between robots and the object to ensure effective and safe collaboration. Due to the tight coupling, any small imbalance or misalignment may cause excessive interaction wrenches by mutual pushing and pulling. The increased wrenches may damage the object or cause the object to slip [11, 12]. Therefore, in object manipulation, the coordination of interaction wrenches is a crucial aspect of collaborative performance.
Generally, the control strategies of MMMS in maintaining desired interaction wrenches can be categorized into centralized, leader-follower, and fully distributed methods. Centralized methods employ a central controller to control the movements and motions of all mobile manipulators [13]. These methods benefit from a global perspective, enabling optimal task allocation and coordination strategies. However, centralized methods suffer from scalability issues as the number of robots increases [14]. The computational burden on the central controller grows, and the system becomes vulnerable to a single point of failure. In recent years, more attention has been given to the latter two methods in research.
The leader-follower method simplifies coordination by assigning one robot as a leader and the rest as followers. Leaders make key decisions, while followers adjust their motions based on the leaders’ states [15, 16]. In [17], a leader-follower structure was used for MMMS to cooperatively transport objects in a constrained workspace. In [18], the coordination was ensured by maintaining the relative pose relationship of the fixed end-effectors between the coupled mobile manipulators. In the simulation or experiment, a robot was chosen as the leader. Similarly, leader-follower control methods can also be found in [19, 20, 21]. While the leader-follower structure reduces control complexity for followers and is practical for real-world implementation, it has notable drawbacks. The system’s performance depends heavily on the leader’s reliability and decision-making, making it vulnerable to single points of failure similar to centralized methods.
Fully distributed methods represent the most advanced approach, where each mobile manipulator makes decisions based on local information and communication with neighboring robots. In [22], a distributed adaptive controller was proposed to achieve synchronization based on the desired load distribution and desired internal force. The effectiveness of their method was validated through experiments. A distributed model predictive control method was presented for MMMS to transport objects with respect to state, control input, and synchronization constraints in [23]. The method simplifies the problem into two independent synchronization tracking control issues for task-space end-effectors and null-space mobile bases. In [24], a chatterless integral sliding mode force-position control was proposed. Each robot was designed to correct collectively the velocity deviation due to the pushing of the other robots. Distributed coordination and cooperation control for networked mobile manipulators over a topology was introduced in [25]. Synchronization control combined with force controls ensured cooperative manipulation and transportation. In [26], a control approach was proposed to guarantee predefined transient and steady-state performance for object trajectory. Load distribution between robots was also considered by using a grasp matrix pseudoinverse.
Table 1: Comparison between distributed control methods.
Torque Control Verification Positioning System
[23, 24, 27, 25, 26] Need Simulation -
[9] No need Simulation -
[22, 28] Need Experiment Need
ours No need Experiment No need
The research on distributed control is summarized in Table 1. It can be seen that most methods require torque control. These methods establish dynamic models to estimate the output wrenches based on the Newton-Euler method or the Lagrangian method [29]. However, in complex multi-degree-of-freedom systems, determining the relevant dynamic parameters presents significant challenges. In these systems, parameters such as mass, inertia, and joint stiffness are critical for control, but they are often difficult to measure or estimate with high precision. Small errors in these parameters can lead to large discrepancies and unwanted performance. Furthermore, joint friction is often ignored or oversimplified in dynamic models. These factors make it difficult to validate their methods in physical systems [23, 24, 27, 25, 26]. In a few research with experimental tests, external positioning systems were used to obtain the motion information of the end-effectors [22, 28].
When an accurate dynamic model is unavailable and torque control is not feasible, the main challenge is how to develop an effective control law with less information. Besides, the design of this control law should be as simple as possible, ensuring that it can be applied in most conditions. Therefore, to address these shortcomings and make MMMS more convenient and applicable in real-world object manipulation, this paper presents a distributed motion control strategy. The robot obtains states through a local sensor and information from neighboring robots. Then, the control inputs for joint velocities are computed to reduce interaction wrenches. In the control process, neither an external positing system nor torque control is required. The stability of the control law has been theoretically proven through the Lyapunov theorem. In the simulations, the effects of graph connectivity and time delays on the control law have been studied. A comparison with other methods has also been made, demonstrating the advantages of the proposed control law. Finally, the control law is validated through physical experiments.
The main contributions of this paper are twofold. First, the proposed control law requires very few conditions, only local force-torque information. Compared to most methods based on dynamic modeling, it does not need torque control, which offers better versatility for different robots, especially for industrial robots without accurate dynamic models. Second, the proposed control law has been validated in physical systems and proven to be robust to communication delays. Most existing studies only provide theoretical proofs and simulations, with no physical validation.
2 Preliminaries
2.1 Kinematics of Mobile Manipulator
@@
A mobile manipulator robot usually consists of a mobile platform and a mounted robotic manipulator. This paper considers an omnidirectional mobile platform and a six-degree-of-freedom manipulator. As shown in Fig. 1, three mobile manipulators cooperatively manipulate objects. $\mathbb{W}$ denotes the frame of the world. $\mathbb{P}_{i}$, $\mathbb{B}_{i}$, and $\mathbb{E}_{i}$ denote the frames of the mobile platform, the base of the manipulator, and the end-effector of the $i$-th mobile manipulator, respectively, $i=1,\cdots,n$. Let $\bf{T}$, $\bf{R}$, $\bm{p}$ represent the relationship between different frames. For example, ${{}^{\mathbb{W}}_{\mathbb{E}_{i}}}{\bf{T}}\in SE(3)$ represents the pose of frame $\mathbb{E}_{i}$ in frame $\mathbb{W}$, ${{}^{\mathbb{W}}_{\mathbb{E}_{i}}}{\bf{R}}\in SO(3)$ represents the orientation of frame $\mathbb{E}_{i}$ in frame $\mathbb{W}$, and ${{}^{\mathbb{W}}_{\mathbb{E}_{i}}}{\bm{p}}\in\mathbb{R}^{3}$ represents the translation of frame $\mathbb{E}_{i}$ in frame $\mathbb{W}$. Based on frame transformations, the following fundamental relationships can be obtained
$^{\mathbb{W}}_{\mathbb{E}_{i}}{\bf{T}}={{}^{\mathbb{W}}_{\mathbb{P}_{i}}}{\bf{T}}\ {{}^{\mathbb{P}_{i}}_{\mathbb{B}_{i}}}{\bf{T}}\ {{}^{\mathbb{B}_{i}}_{\mathbb{E}_{i}}}{\bf{T}}$ (1)
It should be noted that ${{}^{\mathbb{P}_{i}}_{\mathbb{B}_{i}}{\bf{T}}}$ is invariant and solely depends on the location of the manipulator mounted on the mobile platform.
The forward kinematics of the robot is
@@
where $\bm{e}_{i}\in\mathbb{R}^{6}$ is the pose of the end-effector in frame $\mathbb{W}$, $f_{\text{mm}}(\cdot)$ is related to ${{}^{\mathbb{W}}_{\mathbb{E}_{i}}}{\bf{T}}$, and $\bm{q}_{\text{mm},i}\in\mathbb{R}^{9}$ is the configuration of the robot. $\bm{q}_{\text{mm},i}=[\bm{q}_{\text{p},i}^{\text{T}},\bm{q}_{i}^{\text{T}}]^{\text{T}}$, specifically, $\bm{q}_{\text{p},i}=[x,y,\alpha]^{\text{T}}\in\mathbb{R}^{3}$ determining ${{}^{\mathbb{W}}_{\mathbb{P}_{i}}}{\bf{T}}$ denotes the position and orientation of the mobile platform, and $\bm{q}_{i}=[\theta_{1},\theta_{2},\theta_{3},\theta_{4},\theta_{5},\theta_{6}]^{\text{T}}\in\mathbb{R}^{6}$ determining ${{}^{\mathbb{B}_{i}}_{\mathbb{E}_{i}}}{\bf{T}}$ denotes the joint angles of the manipulator. Nonholonomic constraints are not considered in this paper. Thus, differentiating Eq.(2), the velocity relationship is
$\dot{\bm{e}}_{i}=\bm{J}_{\text{mm},i}(\bm{q}_{\text{mm},i})\dot{\bm{q}}_{\text{mm},i}=[\bm{J}_{\text{p},i},\bm{J}_{i}][\bm{q}_{\text{p},i}^{\text{T}},\bm{q}_{i}^{\text{T}}]^{\text{T}}$ (3)
where $\bm{J}_{\text{mm},i}=[\bm{J}_{\text{p},i},\bm{J}_{i}]$ is the Jacobian matrix of the $i$-th robot. $\bm{J}_{\text{p},i}$ and $\bm{J}_{i}$ are the Jacobian matrices of the $i$-th mobile platform and manipulator, respectively.
2.2 Force-Torque Sensor
The manipulator is equipped with a gripper and a force-torque sensor. The data measured by force-torque sensors are usually in their own coordinate systems. Let ${}^{\mathbb{E}_{i}}\bm{h}_{i}=[{{}^{\mathbb{E}_{i}}}\bm{f}_{i}^{\text{T}},{{}^{\mathbb{E}_{i}}}\bm{t}_{i}^{\text{T}}]^{\text{T}}$ represent the data measured by the sensor at the end effector of the $i$-th robot. $\bm{h}_{i}=[\bm{f}_{i}^{\text{T}},\bm{t}_{i}^{\text{T}}]^{\text{T}}$ represents the force and torque exerted on the end-effector of the $i$-th robot in frame $\mathbb{W}$. According to the force transfer relationship, it can be derived
@@ 2.2 Force-Torque Sensor
Before measuring data, it is necessary to perform calibration and gravity compensation for the sensor. In unloaded conditions, the force data consists of
@@
where $\bm{f}_{\text{s}\_0}$ and $\bm{t}_{\text{s}\_0}$ are the drift errors, $\bm{f}_{\text{s}\_g}$ and $\bm{t}_{\text{s}\_g}$ are caused by the gravity of the gripper $mg$. The displacement from the sensor to the gripper’s center of mass is denoted as ${{}^{\mathbb{E}}_{g}}{\bm{p}}$. Parameters $\bm{f}_{\text{s}\_0}$, $\bm{t}_{\text{s}\_0}$, ${{}^{\mathbb{E}}_{g}}{\bm{p}}$, and $mg$, which need further calibration, are linearly independent. Therefore, these parameters can be calculated using the least squares method after obtaining enough sensor data in unloaded conditions [30].
@@
$\bm{G}$ $=\left[\begin{matrix}I_{3}&0_{3}&\cdots&I_{3}&0_{3}\\ [\bm{r}_{1}]_{\times}&I_{3}&\cdots&[\bm{r}_{n}]_{\times}&I_{3}\end{matrix}\right]$
where $\bm{G}_{6\times 6n}$ is the grasp matrix, which is only related to the relative poses of the grasping points and the object’s center of mass [31, 32, 22], as the $\bm{r}_{1},\bm{r}_{2},\bm{r}_{3}$ shown in Fig. 1. $[\bm{r}_{i}]_{\times}$ represents the skew-symmetric matrix of $\bm{r}_{i}$.
2.3 Graph Theory
The communication structure of a system with $n$ robots can be typically represented on graphs. Let $\mathcal{G=(V,E)}$ denote a graph, where $\mathcal{V}=\left\{v_{i}\right\}$ is the vertex set and =$\mathcal{E}=\left\{a_{ij}\right\}$ is the edge set. $a_{ij}=1$ if the $i$-th robot can receive information from the $j$-th robot, otherwise $a_{ij}=0$. The adjacency matrix of the graph is $A=(a_{ij})\in\mathbb{R}^{n\times n}$. The degree matrix of the graph is $D=\text{diag}\left\{\sum_{j=1}^{n}a_{ij}\right\}\in\mathbb{R}^{n\times n}$. The Laplacian matrix of the graph is $L=D-A$. In this paper, the directed graph is considered, which denotes $a_{ij}\neq a_{ij}$.
3 The Proposed Control Method
3.1 Problem Statement and Control Law